OpenStudy (anonymous):

what is the rule for finding the 100th term in the sequence 0,2,4,8,16

OpenStudy (anonymous):

any help will be appreciated. thank you very much!

OpenStudy (erinweeks):

mult. by 2

OpenStudy (anonymous):

By rule i mean the equation

OpenStudy (anonymous):

So the rule for finding the nth term

OpenStudy (anonymous):

Are you sure it isn't 1,2,4,8,16

OpenStudy (anonymous):

\[ a_n = a_1\cdot \left(\frac {a_2}{a_1}\right)^{n-1} \]

OpenStudy (anonymous):

This is the formula for an geometric sequence given the first two terms.

OpenStudy (anonymous):

It's not arithmetic.

OpenStudy (anonymous):

It's geometric if the first term is corrected.

OpenStudy (anonymous):

Yes, to be geometric, the first term must be removed.